# Measurability in the Fundamental Theorem of Statistical Learning

### v2

@@ Abstract
Math Subject Classification (2020): Primary 68T05, 03C64; Secondary 28A05, 28A20, 03C98, 68T27, 12J15. Key Words: PAC Learning, VC Dimension, Uniform Convergence, Measurability, Borel Sets, NIP, o-Minimality. Corresponding Author: Laura Wirth, laura@uni-konstanz.de.
@@ 1 Introduction
In 1984, Valiant [32] introduced a formal model of learning based on randomly generated examples, which is referred to as Probably Approximately Correct (PAC) Learning. Conceptually, a class of (indicator) functions is PAC learnable if for any finitely many randomly generated examples there exists a suitable generalization function within the class. Blumer, Ehrenfeucht, Haussler and Warmuth [5] established in 1989 that an indicator function class is PAC learnable if and only if its VC dimension is finite. Their result is commonly referred to as the Fundamental Theorem of Statistical Learning. Both the argument in [5] and all following proofs of this theorem in the literature rely on “certain broadly applicable measurability conditions” (Ackerman, Asilis, Di, Freer and Tristan [1, page 1]). The VC dimension – introduced by Vapnik and Chervonenkis [^1] Due to different transliteration from the Cyrillic, various spellings of this name are common in Latin script. [33] in 1968 – is a parameter for measuring the combinatorial size of an indicator function class. Laying the basis for the Fundamental Theorem of Statistical Learning, Vapnik and Chervonenkis [34] proved in 1971 that finiteness of the VC dimension is a necessary and sufficient condition for uniform convergence of empirical means. In 1992, Laskowski [18] revealed a striking analogy between NIP formulas and classes having finite VC dimension. Here, NIP stands for ‘not the independence property’ (for details see Poizat [24, § 12.4]). This triangular relationship of the notions NIP, VC dimension and PAC learnability inspired a fruitful exchange between Model Theory and Statistical Learning Theory, as “connections between the VC theory and NIP have developed extensively with important notions from VC-theory adapted to the model-theoretic setting and vice versa” (Chase and Freitag [8, page 323]). For instance, Steinhorn [31, page 27] points out that “a definable family of definable sets $\mathcal{C}$ in an o-minimal structure is PAC learnable”. For drawing such connections special care has to be taken, as the Fundamental Theorem of Statistical Learning applies only to classes satisfying certain measurability conditions. In [5], a precise analysis of the measurability requirements is provided. The authors introduce the notion of well-behaved classes [^2] This is a measure-theoretic property that is developed in [5, Appendix A1]. We introduce a similar property in Section 3., to which the Fundamental Theorem of Statistical Learning applies. Although in [5, page 935] this property is described as “benign”, it requires “the measurability of complex sets of samples” (Ben-David, Benedek and Mansour [3, page 13]). Thus, the well-behavedness assumption constitutes a “complicated measurability constraint” (Ben-David, Itai and Kushilevitz [4, page 248]).
@@
The underlying PAC learning model considered in [5, § 2] relies on deterministic labels. More precisely, the samples on which the learning is based are generated according to distributions on the instance space and are equipped with labels that are generated according to a target function. In contrast, following more recent literature such as Shalev-Shwartz and Ben-David [27, § 3.2.1], we work with a model of PAC learning that is agnostic in various respects. This model omits target functions and is instead based on working with an arbitrary set of joint distributions on the sample space. Thus, the labels are not deterministic but instead also generated randomly. The agnostic PAC model therefore provides a more general and realistic description of formal learning (cf. Vidyasagar [36, § 3.3.1]). In particular, it is able to deal with noisy and corrupted data (cf. von Luxburg and Schölkopf [19, page 654 f.]). For further details see Remark 2.2, cf. also [5, Appendix A.3].
Several expositions in Statistical Learning acknowledge that measurability subtleties demand a certain amount of attention (see e.g. Haussler [14, page 190], Mendelson and Smola [21, page 5], Pestov [22, page 1141] and Vidyasagar [36, page 56]). However, the available literature so far lacks a proof of the Fundamental Theorem of Statistical Learning that rigorously takes into account the measurability requirements of agnostic PAC learning. In this article, we fill this gap by presenting a self-contained proof that addresses the involved measurability issues in a detailed and precise manner.
Our initial motivation for the present work was to investigate further applications of the Fundamental Theorem of Statistical Learning in the model-theoretic context of linearly ordered structures. To this end, we extract the measure-theoretic prerequisites in order to build a foundation upon which future model-theoretic results can be precisely formulated.
@@
Throughout this paper, we focus on binary classification problems, as these are well-studied within Supervised Learning, a subbranch of Machine Learning. In this context, (supervised) learning refers to the process of inferring a general rule on how to classify certain objects by observing examples.
@@
Our paper is structured as follows. In Section 2 we describe our learning framework and gather basic preliminaries from Statistical Learning Theory. In particular, we introduce the notions that are fundamental for the results we establish in Section 3: (agnostic) PAC Learning, Uniform Convergence Property and VC Dimension. Special emphasis is put on measure-theoretic aspects of these concepts. In Section 3 we present a sound statement as well as a detailed and self-contained proof of the Fundamental Theorem of Statistical Learning (Theorem 3.3). Its result relies on some measurability assumptions that are described in Definition 3.2. The proof comprises the verification of several implications (see Theorem 3.8, Theorem 3.9, Proposition 3.10 and Theorem 3.11). The technical heart is to prove that any hypothesis space with finite VC dimension has the uniform convergence property, provided certain measurability requirements are met (see Theorem 3.6 and Theorem 3.8). Section 4 initiates an examination of applications of the Fundamental Theorem of Statistical Learning in Model Theory. We introduce definable hypothesis spaces over first-order structures (Definition 4.1), which always have finite VC dimension if the underlying structure has NIP (Proposition 4.5). Our main theorem (Theorem 4.7) establishes sufficient conditions for the PAC learnability of hypothesis spaces defined over o-minimal expansions of the reals. This class of hypothesis spaces encompasses artificial neural networks whose activation functions are definable over the real exponential field (Example 4.2).We complete this work by gathering open questions in Section 5.
@@
Acknowledgements: We thank Matthias C. Caro, Salma Kuhlmann andTobias Sutter for several helpful discussions and for valuable advice on relevant literature. The second author would also like to extend special thanks to Salma Kuhlmann for the supervision of her doctoral project, which this work is part of. Additionally, the second author is grateful to David Pollard, Itay Kaplan, Markus Kunze, Tim Seitz and Patrick Michalski for pointing out various challenges related to measurability, clarifying some subtleties, and suggesting potential solutions during personal exchanges.
@@ 2 Framework and Preliminaries
We denote by $\mathbb{N}$ the set of positive natural numbers and we set $\mathbb{N}_{0}=\mathbb{N}\,\dot{\cup}\,\{0\}$. Given $m\in\mathbb{N}$, we denote by $[m]$ and $[m]_{0}$ the sets $\{1,\dots,m\}$ and $\{0,\dots,m\}$, respectively. Often we would like to emphasize that some object is a tuple and we then use underlined letters such as $\underline{z}$. The power set of a set $A$ is denoted by $\mathcal{P}(A)$, and the set of finite subsets of $A$ by $\mathcal{P}_{\operatorname{fin}}(A)$. We denote by $f\mathord{\upharpoonright}_{A}$ the restriction of a function $f$ to a subset $A$ of its domain. The indicator function $\mathds{1}_{A}$ of a set $A$ maps any element of $A$ to $1$ and everything else to 00. Given a measurable space $(\Omega,\Sigma)$ and an element $\omega\in\Omega$, we denote by $\delta_{\omega}$ the Dirac measure defined by
@@
    The instance space $\mathcal{X}$ contains the objects to be classified into positive and negative examples using the binary labels 00 and $1$.
@@
that are drawn independently at random according to some probability distribution $\mathbb{D}\in\mathcal{D}$. Thus, the multi-sample $\underline{z}=((x_{1},y_{1}),\dots,(x_{m},y_{m}))\in\mathcal{Z}^{m}$ is drawn according to $\mathbb{D}^{m}$, where $\mathbb{D}^{m}$ denotes the product measure $\bigotimes_{i=1}^{m}\mathbb{D},$ which is defined as usual on the product $\sigma$–algebra [^5] The product $\sigma$–algebra $\Sigma_{\mathcal{Z}}^{m}=\bigotimes_{i=1}^{m}\Sigma_{\mathcal{Z}}$ on $\mathcal{Z}^{m}$ is the smallest $\sigma$–algebra on $\mathcal{Z}^{m}$ containing all sets of the form ×$\bigtimes\limits_{i\in[m]}\!C_{i}$, where $C_{i}\in\Sigma_{\mathcal{Z}}$ for $i\in[m]$. $\Sigma_{\mathcal{Z}}^{m}$.
@@
As we consider joint distributions $\mathbb{D}$ on the sample space $\mathcal{Z}$, it is possibly the case that for an instance $x\in\mathcal{X}$ both $(x,0)$ and $(x,1)$ have positive probability. In other words, the labels of instances are not deterministic or provided by a specific target function but are random themselves. This also allows us to deal with “noisy labels or corrupted data”. The underlying learning paradigm is sometimes referred to as agnostic (see [27, § 3.2.1]) or model-free (see [36, § 3.3], respectively). In contrast, Blumer, Ehrenfeucht, Haussler and Warmuth [5, § 2] work with a learning model that only involves distributions on the instance space $\mathcal{X}$. Thus, the instances $x_{1},\dots,x_{m}\in\mathcal{X}$ are drawn independently at random according to some probability distribution $\mathbb{P}$ on a $\sigma$–algebra on $\mathcal{X}$. The labels $y_{1},\dots,y_{m}$ are then generated according to a fixed target function $t\in\{0,1\}^{\mathcal{X}}$.
@@
where :×$\ell\colon\mathcal{H}\times\mathcal{Z}\to\{0,1\}$, referred to as loss function, is defined by
@@
$\mathbb{E}_{\underline{z}\sim\mathbb{D}^{m}}[\hat{\operatorname{er}}_{\underline{z}}(h)]=\frac{1}{m}\sum\limits_{i=1}^{m}\mathbb{E}_{z\sim\mathbb{D}}[\ell(h,z)]=\mathbb{E}_{z\sim\mathbb{D}}[\ell(h,z)]=\operatorname{er}_{\mathbb{D}}(h).\vspace{-0.7em}$
Following [27, Definition 4.3] and [36, § 3.1.2], we now introduce the uniform convergence property, which guarantees that sample errors are “good” estimates for true errors.
@@
$\underline{z}$ $\mapsto\sup\limits_{h\in\mathcal{H}}\big{|}\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big{|}$
@@
for $m\in\mathbb{N}$. The following result allows us to polynomially bound the values of the growth function using the VC dimension. It was established independently by Sauer [26, Theorem 1], Shelah [29], Vapnik and Chervonenkis [34, § 2], with motivation in different areas. As Shelah [29, page 254] also mentions that Perles was involved in the proof, it is sometimes called the Sauer–Shelah–Perles Lemma. Others refer to the result simply as Sauer’s Lemma.
@@
We finally specify the formal model of learning that we work with: Probably Approximately Correct (PAC) Learning. This concept, first introduced by Valiant [32], was modified and extended over time to also capture the agnostic setup that we deal with. The definition we use is a refined version of [2, Definition 2.1], [27, Definition 3.3] and [36, Definition 3.5], further inspired by [5, § 2].
@@
and $\text{and }\mathbb{D}^{m}(C)\geq 1-\delta.$
@@
Note that the set $\{\underline{z}\in\mathcal{Z}^{m}\mid\operatorname{er}_{\mathbb{D}}(\mathcal{A}(\underline{z}))-\operatorname{opt}_{\mathbb{D}}(\mathcal{H})\leq\varepsilon\}$ is not necessarily contained in $\Sigma_{\mathcal{Z}}^{m}$ (see Example A.12). [^7] For this reason, our notion of PAC learnability employs the computation of the inner measure, which can be computed also for sets that are not measurable. Further, we point out that in our model, PAC learnability of hypothesis spaces can be expressed with respect to any set $\mathcal{D}$ of distributions defined on the measurable space $(\mathcal{Z},\Sigma_{\mathcal{Z}})$. In the literature, however, PAC learnability is often defined in terms of the class of all (suitable) distributions (see e.g. [2, Definition 2.1] [27, Definition 3.3]).

### v3

@@ Abstract
Math Subject Classification (2020): Primary 68T05, 03C64; Secondary 28A05, 28A20, 03C98, 68T27, 12J15. Key Words: PAC Learning, VC Dimension, Uniform Convergence, Measurability, NIP, o-Minimality. Corresponding Author: Laura Wirth, laura@uni-konstanz.de.
@@ 1 Introduction
In 1984, Valiant [33] introduced a formal model of learning based on randomly generated examples, which is referred to as Probably Approximately Correct (PAC) Learning. Conceptually, a class of (indicator) functions is PAC learnable if for any finitely many randomly generated examples there exists a suitable generalization function within the class. Blumer, Ehrenfeucht, Haussler and Warmuth [5] established in 1989 that an indicator function class is PAC learnable if and only if its VC dimension is finite. Their result is commonly referred to as the Fundamental Theorem of Statistical Learning. Both the argument in [5] and all following proofs of this theorem in the literature rely on “certain broadly applicable measurability conditions” (Ackerman, Asilis, Di, Freer and Tristan [1, page 1]). The VC dimension – introduced by Vapnik and Chervonenkis [^1] Due to different transliteration from the Cyrillic, various spellings of this name are common in Latin script. [34] in 1968 – is a parameter for measuring the combinatorial size of an indicator function class. Laying the basis for the Fundamental Theorem of Statistical Learning, Vapnik and Chervonenkis [35] proved in 1971 that finiteness of the VC dimension is a necessary and sufficient condition for uniform convergence of empirical means. In 1992, Laskowski [18] revealed a striking analogy between NIP formulas and classes having finite VC dimension. Here, NIP stands for ‘not the independence property’ (for details see Poizat [25, § 12.4]). This triangular relationship of the notions NIP, VC dimension and PAC learnability inspired a fruitful exchange between Model Theory and Statistical Learning Theory, as “connections between the VC theory and NIP have developed extensively with important notions from VC-theory adapted to the model-theoretic setting and vice versa” (Chase and Freitag [8, page 323]). For instance, Steinhorn [32, page 27] points out that “a definable family of definable sets $\mathcal{C}$ in an o-minimal structure is PAC learnable”. For drawing such connections special care has to be taken, as the Fundamental Theorem of Statistical Learning applies only to classes satisfying certain measurability conditions. In [5], a precise analysis of the measurability requirements is provided. The authors introduce the notion of well-behaved classes [^2] This is a measure-theoretic property that is developed in [5, Appendix A1]. We introduce a similar property in Section 3., to which the Fundamental Theorem of Statistical Learning applies. Although in [5, page 935] this property is described as “benign”, it requires “the measurability of complex sets of samples” (Ben-David, Benedek and Mansour [3, page 13]). Thus, the well-behavedness assumption constitutes a “complicated measurability constraint” (Ben-David, Itai and Kushilevitz [4, page 248]).
@@
The underlying PAC learning model considered in [5, § 2] relies on deterministic labels. More precisely, the samples on which the learning is based are generated according to distributions on the instance space and are equipped with labels that are generated according to a target function. In contrast, following more recent literature such as Shalev-Shwartz and Ben-David [28, § 3.2.1], we work with a model of PAC learning that is agnostic in various respects. This model omits target functions and is instead based on working with an arbitrary set of joint distributions on the sample space. Thus, the labels are not deterministic but instead also generated randomly. The agnostic PAC model therefore provides a more general and realistic description of formal learning (cf. Vidyasagar [36, § 3.3.1]). In particular, it is able to deal with noisy and corrupted data (cf. von Luxburg and Schölkopf [20, page 654 f.]). For further details see Remark 2.2, cf. also [5, Appendix A.3].
Several expositions in Statistical Learning acknowledge that measurability subtleties demand a certain amount of attention (see e.g. Haussler [14, page 190], Mendelson and Smola [22, page 5], Pestov [23, page 1141] and Vidyasagar [36, page 56]). Some authors assume that the function class under consideration is permissible [^3] This is a measure-theoretic property that is developed in Pollard [26, Appendix C]. (e.g. Haussler [15, § 9.2] and Lee [19, § 2.3]) in order to prove according results. From a purely theoretical point of view, this assumption confines consideration to function classes with a rather specific structure. In practice, most function classes arising in applications satisfy this condition.
Our initial motivation for this work was to explore theoretical applications of the Fundamental Theorem of Statistical Learning, particularly in the model-theoretic context of linearly ordered structures. However, the available literature so far lacks a systematic analysis of the minimal measurability requirements of the Fundamental Theorem of Statistical Learning for agnostic PAC learning. In this article, we fill this gap by identifying these requirements, and thereby describing the most general measure-theoretic framework in which the result and its proof remain valid. More precisely, we present a self-contained proof that addresses the relevant measurability issues in a detailed and precise manner. In doing so, we build a foundation upon which future learnability results in the context of Model Theory can be precisely formulated.
@@
Throughout this paper, we concentrate on binary classification problems, as these are well-studied within Supervised Learning, a subbranch of Machine Learning. In this context, (supervised) learning refers to the process of inferring a general rule on how to classify certain objects by observing examples. Our focus lies on establishing qualitative learnability results, and we therefore do not address quantitative aspects such as explicit sample complexity bounds.
@@
Our paper is structured as follows. In Section 2 we describe our learning framework and gather basic preliminaries from Statistical Learning Theory. In particular, we introduce the notions that are fundamental for the results we establish in Section 3: (agnostic) PAC Learning, Uniform Convergence Property and VC Dimension. Special emphasis is put on measure-theoretic aspects of these concepts. In Section 3 we present a sound statement as well as a detailed and self-contained proof of the Fundamental Theorem of Statistical Learning (Theorem 3.3). Its result relies on some measurability assumptions that are described in Definition 3.2. The proof comprises the verification of several implications (see Theorem 3.8, Theorem 3.9, Proposition 3.10 and Theorem 3.11). A key step is to prove that any hypothesis space with finite VC dimension has the uniform convergence property, provided certain measurability requirements are met (see Theorem 3.6 and Theorem 3.8). This is the part of the argument that demands most of the technical work and is precisely where measurability assumptions are crucial. Section 4 initiates an examination of applications of the Fundamental Theorem of Statistical Learning in Model Theory. We introduce definable hypothesis spaces over first-order structures (Definition 4.1), which always have finite VC dimension if the underlying structure has NIP (Proposition 4.5). Our main theorem (Theorem 4.7) establishes sufficient conditions for the PAC learnability of hypothesis spaces defined over o-minimal expansions of the reals. This class of hypothesis spaces encompasses artificial neural networks whose activation functions are definable over the real exponential field (Example 4.2).We complete this work by gathering open questions in Section 5.
@@
Acknowledgements: We thank Matthias C. Caro, Salma Kuhlmann andTobias Sutter for several helpful discussions and for valuable advice on relevant literature. Moreover, we thank anonymous referees for providing several helpful comments and references. The second author would also like to extend special thanks to Salma Kuhlmann for the supervision of her doctoral project, which this work is part of. Additionally, the second author is grateful to David Pollard, Itay Kaplan, Markus Kunze, Tim Seitz and Patrick Michalski for pointing out various challenges related to measurability, clarifying some subtleties, and suggesting potential solutions during personal exchanges.
@@ 2 Framework and Preliminaries
We denote by $\mathbb{N}$ the set of positive natural numbers and we set $\mathbb{N}_{0}=\mathbb{N}\,\dot{\cup}\,\{0\}$. Given $m\in\mathbb{N}$, we denote by $[m]$ and $m]_{0}$ the sets $\{1,\dots,m\}$ and $\{0,\dots,m\}$, respectively. Often we would like to emphasize that some object is a tuple and we then use underlined letters such as $\underline{z}$. The power set of a set $A$ is denoted by $\mathcal{P}(A)$, and the set of finite subsets of $A$ by $\mathcal{P}_{\operatorname{fin}}(A)$. We denote by $f\mathord{\upharpoonright}_{A}$ the restriction of a function $f$ to a subset $A$ of its domain. The indicator function $\mathds{1}_{A}$ of a set $A$ maps any element of $A$ to $1$ and everything else to 0. Given a measurable space $(\Omega,\Sigma)$ and an element $\omega\in\Omega$, we denote by $\delta_{\omega}$ the Dirac measure defined by
@@
    The instance space $\mathcal{X}$ contains the objects to be classified into positive and negative examples using the binary labels 0 and $1$.
@@
that are drawn independently at random according to some probability distribution $\mathbb{D}\in\mathcal{D}$. Thus, the multi-sample $\underline{z}=((x_{1},y_{1}),\dots,(x_{m},y_{m}))\in\mathcal{Z}^{m}$ is drawn according to $\mathbb{D}^{m}$, where $\mathbb{D}^{m}$ denotes the product measure $\bigotimes_{i=1}^{m}\mathbb{D},$ which is defined as usual on the product $\sigma$–algebra [^6] The product $\sigma$–algebra $\Sigma_{\mathcal{Z}}^{m}=\bigotimes_{i=1}^{m}\Sigma_{\mathcal{Z}}$ on $\mathcal{Z}^{m}$ is the smallest $\sigma$–algebra on $\mathcal{Z}^{m}$ containing all sets of the form $\bigtimes\limits_{i\in[m]}\!C_{i}$, where $C_{i}\in\Sigma_{\mathcal{Z}}$ for $i\in[m]$. $\Sigma_{\mathcal{Z}}^{m}$.
@@
As we consider joint distributions $\mathbb{D}$ on the sample space $\mathcal{Z}$, it is possibly the case that for an instance $x\in\mathcal{X}$ both $(x,0)$ and $(x,1)$ have positive probability. In other words, the labels of instances are not deterministic or provided by a specific target function but are random themselves. This also allows us to deal with “noisy labels or corrupted data”. The underlying learning paradigm is sometimes referred to as agnostic (see [28, § 3.2.1]) or model-free (see [36, § 3.3], respectively). In contrast, Blumer, Ehrenfeucht, Haussler and Warmuth [5, § 2] work with a learning model that only involves distributions on the instance space $\mathcal{X}$. Thus, the instances $x_{1},\dots,x_{m}\in\mathcal{X}$ are drawn independently at random according to some probability distribution $\mathbb{P}$ on a $\sigma$–algebra on $\mathcal{X}$. The labels $y_{1},\dots,y_{m}$ are then generated according to a fixed target function $t\in\{0,1\}^{\mathcal{X}}$.
@@
where $\ell\colon\mathcal{H}\times\mathcal{Z}\to\{0,1\}$, referred to as loss function, is defined by
@@
$\mathbb{E}_{\underline{z}\sim\mathbb{D}^{m}}[\hat{\operatorname{er}}_{\underline{z}}(h)]=\frac{1}{m}\sum\limits_{i=1}^{m}\mathbb{E}_{z\sim\mathbb{D}}[\ell(h,z)]=\mathbb{E}_{z\sim\mathbb{D}}[\ell(h,z)]=\operatorname{er}_{\mathbb{D}}(h).\vskip-6.99997pt$
Following [28, Definition 4.3] and [36, § 3.1.2], we now introduce the uniform convergence property, which guarantees that sample errors are “good” estimates for true errors.
@@
$\underline{z}$ $\mapsto\sup\limits_{h\in\mathcal{H}}\big|\hskip-1.25pt\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big|$
@@
for $m\in\mathbb{N}$. The following result allows us to polynomially bound the values of the growth function using the VC dimension. It was established independently by Sauer [27, Theorem 1], Shelah [30], Vapnik and Chervonenkis [35, § 2], with motivation in different areas. As Shelah [30, page 254] also mentions that Perles was involved in the proof, it is sometimes called the Sauer–Shelah–Perles Lemma. Others refer to the result simply as Sauer’s Lemma.
@@
We finally specify the formal model of learning that we work with: Probably Approximately Correct (PAC) Learning. This concept, first introduced by Valiant [33], was modified and extended over time to also capture the agnostic setup that we deal with. The definition we use is a refined version of [2, Definition 2.1], [28, Definition 3.3] and [36, Definition 3.5], further inspired by [5, § 2].
@@
$\text{and }\mathbb{D}^{m}(C)\geq 1-\delta.$
@@
Note that the set $\{\underline{z}\in\mathcal{Z}^{m}\mid\operatorname{er}_{\mathbb{D}}(\mathcal{A}(\underline{z}))-\operatorname{opt}_{\mathbb{D}}(\mathcal{H})\leq\varepsilon\}$ is not necessarily contained in $\Sigma_{\mathcal{Z}}^{m}$ (see Example A.12). [^8] For this reason, our notion of PAC learnability employs the computation of the inner measure, which can be computed also for sets that are not measurable. Further, we point out that in our model, PAC learnability of hypothesis spaces can be expressed with respect to any set $\mathcal{D}$ of distributions defined on the measurable space $(\mathcal{Z},\Sigma_{\mathcal{Z}})$. In the literature, however, PAC learnability is often defined in terms of the class of all (suitable) distributions (see e.g. [2, Definition 2.1] [28, Definition 3.3]).